OpenStudy (anonymous):

solve using the substitution method y = 5 - x 2x + 3y = 12

OpenStudy (johnweldon1993):

Well as you see, that first equation...tells you what 'y' equals... in your second equation...now that you know what 'y' equals...you can substitute that in for every 'y' you see y = 5 - x 2x + 3y = 12 after substitution 2x + 3(5-x) = 12 now solve for 'x'